OpenStudy (anonymous):

if student has three test marks of 73%, 82%, and 65% and wants an average of 80% over four tests, what must be her mark on the next test be?

hero (hero):

(73+82+65+x)/4 = 80

OpenStudy (anonymous):

100!!!

OpenStudy (anonymous):

100%

hero (hero):

73+82+65+x = 320 220 + x = 320 x = 100

OpenStudy (lalaly):

iza ma fhemti bfahmkk kf :D

OpenStudy (anonymous):

shukran fhmta

OpenStudy (lalaly):

ok=)

hero (hero):

What language is that?

OpenStudy (anonymous):

what language is that?

OpenStudy (anonymous):

LOLOL

OpenStudy (lalaly):

lol arabic

hero (hero):

Tomas, don't repeat after me :P

OpenStudy (anonymous):

arabic

OpenStudy (anonymous):

مرحبا انها لغة غريبة حقا

OpenStudy (lalaly):

hehehe :P

OpenStudy (anonymous):

hahahhha

hero (hero):

rld found me again

OpenStudy (anonymous):

Oh u were hiding?

hero (hero):

Maybe

OpenStudy (anonymous):

That is what i thought :P